OpenStudy (anonymous):

The formula that would proof this question true or false is? If a force of 10 N is applied to a 4.5 kg object, then the object will accelerate at 2.22 m/s2.

OpenStudy (anonymous):

yes, it's true

OpenStudy (ganpat):

F = Ma.. Newtons second law..

OpenStudy (anonymous):

F=M*A 10=4.5*2.22 which is ~10 so yes Newtons 2nd law proves this question

OpenStudy (anonymous):

F=M*a